OpenStudy (anonymous):

i need help figuring this out use matrices abc and d to find each scalar product and sum or difference if possible if a operation is not defined lable it undefined

OpenStudy (anonymous):

a=( 6 1 0 8/-4 3 7 11) b=(1 3/ -2 4) c=(-2 1/ 4 0/2 2/1 1) d=(5 -2/ 3 6)

OpenStudy (anonymous):

12. b-2a

OpenStudy (anonymous):

\[a=\left[\begin{matrix}6 & 1 &0 & 8 \\ -4 & 3 & 7 & 11\end{matrix}\right]\]\[b=\left[\begin{matrix}1 & 3 \\ -2 & 4\end{matrix}\right]\], right?

OpenStudy (anonymous):

yes exactly

OpenStudy (anonymous):

you can't do b-2a... \[2a=\left[\begin{matrix}12 & 2 & 0 & 16 \\ -8 & 6 & 14 & 22\end{matrix}\right]\]but its dimension is 2x4... b has dimension 2x2. can't add/subtract matrices of differeing dimension

OpenStudy (anonymous):

so it would be undefined?

OpenStudy (anonymous):

thank you and if you dont mind i have 4 more similar to this one if you would like to help me i would greatly apreciate it

OpenStudy (anonymous):

what do you think? add/subtaction of matrices is defined by adding/subtracting corresponding entries. if the matrices have differing dimensions, they won't have all corresponding entries.

OpenStudy (anonymous):

fire away

OpenStudy (anonymous):

i think its undefined because you cant solve itand thank you

OpenStudy (anonymous):

13.AB 14.BA 15.AC-BD 16.4B-3D

OpenStudy (anonymous):

do you know matrix arithmetic at all?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

i suggest you check out these sites and come back here if you have remaining questions. http://www.purplemath.com/modules/mtrxadd.htm http://www.purplemath.com/modules/mtrxmult.htm

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

@pgpilot326 i have a question the site for multiplication says multiply each row by the colums so in my case of AB i would do 6*1 6*2 and so on?

OpenStudy (anonymous):

nvmi get it now

OpenStudy (anonymous):

remember, to do multiplication of matrices, the number of columns of the left matrix must equal the number of rows for the right matrix. so for AB we have A(2,4) or A has dimension 2x4 meaning A has 2 rows and 4 columns. B(2,2) or B has dimension 2x2 (2 rows and 2 columns) SInce A is on the left, and B is on the right, does the number of columns of A match the number of rows of B? 4 <> 2 so no. Thus, we cannot multiply B by A on the left, i.e., cannot do AB.

OpenStudy (anonymous):

oh so i did jus did all this work for nun i thought i had it

OpenStudy (anonymous):

.AB=(6*1=6+1*-2=-2+0*1=0+8*-2=-16) (6*3=18+1*4=4+0*3=0+8*4=32) (-4*1=-4+3*-2=-6+7*1=7+11*-2=-22) (-4*3=-12+3*4=12+7*3=21+11*4=44) (6+-2+0+-16) (18+4+0+32) /(-4+-6+7+-22) (-12+12+21+44) (-12 54/-25 65)

OpenStudy (anonymous):

@pgpilot326

OpenStudy (anonymous):

BA=(1*6=6+3*1=3+1*0=0+3*8=24) (1*-4=-4+3*3=9+1*7=7+3*11) (-2*6=-12+4*1=4+-2*0=0+4*8=32) (-2*-4=8+4*3=12+-2*7=-14+4*11=44) (6+3+0+24) (-4+9+7+33) /(-12+4+0+32) (8+12+-14+44) (33 45/24 45)